OpenStudy (lilai3):

Hey! I <333333 fruity fruits tooo!!!!! <3333333 A farmer sells a bushel of soybeans for $5 and a bushel of corn for $3. If he hopes to earn $164 and plans to sell 40 bushels in all, how many bushels of soybeans does he need to sell?

OpenStudy (anonymous):

\[(5x+3y)=164 \]

OpenStudy (anonymous):

and \[x+y=40\]

OpenStudy (anonymous):

seewht you can do with those first

OpenStudy (lilai3):

uhm...idk

OpenStudy (anonymous):

all right this needs some trial and error work first write down possible combinations to get to forty that seem reasonable.

OpenStudy (anonymous):

then plug in these numbers until you get an answer that fits.

OpenStudy (lilai3):

uhmm....5(40-y)+ 3y=164.... ???? idk if it's right most likely it's gonna be wrong :P

OpenStudy (anonymous):

try it.

OpenStudy (anonymous):

the worst possible outcome is that you are wrong. its no big deal. just plug your answer it into the original two i gave you to check if you are right.

OpenStudy (lilai3):

then i don't get it at all then

OpenStudy (anonymous):

i hope im not making things worse.

OpenStudy (lilai3):

don worry ur not :DDD

OpenStudy (accessdenied):

Or you could solve the system using Algebra 5x + 3y = 164 x + y = 40 This one is pretty easy to solve using substitution or elimination. I think solving it using the Algebra would be a little more direct than trial than error, at least.

OpenStudy (lilai3):

that's y its called help and thank you @accessdenied!

OpenStudy (lilai3):

uhm....?/?????

OpenStudy (accessdenied):

Well, if we wanted to use substitution, we could solve for a variable in the second equation, and then substitute it back into the first to solve since it'll be one-variable. x + y = 40 - y -y x = 40 - y Which we can plug back into the first: 5(x) + 3y = 164 Replace 'x' with above 5(40 - y) + 3y = 164 then from there we could solve for the y-value, and plug it back into the first equation.

OpenStudy (accessdenied):

plug the y-value back into the second equation, rather. (I am referring to the 1st equation as 5x + 3y = 164 and 2nd is x + y = 40)

OpenStudy (lilai3):

thank you i understand now thank you SO VERY MUCH!!!!!!!!!!!!!!!!!!!!!

OpenStudy (accessdenied):

You're welcome. :)

OpenStudy (lilai3):

:DDD hope you have a great day

OpenStudy (accessdenied):

Thank you. To you as well. :)

OpenStudy (lilai3):

wait i haf a queston: uhm....what is the equation thtat you plug in and what did you geT?

OpenStudy (accessdenied):

5(40 - y) + 3y = 164 200 - 5y + 3y = 164 200 - 2y = 164 100 - y = 82 -y = -18 y = 18 ---> x + y = 40 x + 18 = 40 x = 22

OpenStudy (lilai3):

omfg now for REAL i understand

OpenStudy (accessdenied):

Hmm. :P Good!

OpenStudy (lilai3):

whew i wub ur gelp, bro!